OpenStudy (anonymous):

If xy = 2 and x^2 + y^2 = 5, then x/y + y/x =

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

x/y +y/x = x^2 + y^2 /xy = 5 / 2 (From given equaions)

OpenStudy (anonymous):

thanks thanks ..

OpenStudy (lgbasallote):

if you like some solution... xy = 2 so.. x = 2/y x^2 + y^2 = 5 so... x = sqrt (5 - y^2) x = x so... 2/y = sqrt 95-y^2) square both sides... 4/y^2 = 5 - y^2 multiply everything by y^2 to bring it on top... 4 = 5y^2 - y^4 y^4 - 5y^2 + 4 = 0 let a = y^2 a^2 - 5a + 4 = 0 (a-1)(a-4) = 0 a = 1; 4 y^2 = 1, 4 y = 1, - 1, 2, -2 xy = 2 if y = 1..x = 2 y = -1 x = -2 y = 2..x = 4 y = -2 x = -4 so... x/y + y/x = 2/1 + 1/2 = 4/2 = 1/2 = 5/2 (it is the same for all values of x and y)

OpenStudy (lgbasallote):

that las one was 4/2 + 1/2 = 5/2